a lot is 75m by 50m. A house 24m by 10m is built on the lot. How much area is left over for the lawn
1
Expert's answer
2011-10-11T10:47:24-0400
Lot area is L = 75*50 = 3750 square meters;
house area is H = 24*10 = 240 square meters;
then the rest area for the lawn is L - H = 3750 - 240 = 3510 square meters.